OpenStudy (anonymous):

After the first term in a sequence of positive integers, the ratio of each term to the term immediately preceding it is 3 to 1. What is the ratio of the 105th in this sequence to the 103rd term?

OpenStudy (anonymous):

Translate the mathematical terminology in the sentences into an algebraic identity relating the terms in the series.

OpenStudy (anonymous):

is it 9 to 1?

OpenStudy (anonymous):

$$\frac{a_n}{a_{n+1}}=3$$ For any natural number n

OpenStudy (anonymous):

There asking for $$\frac{a_{105}}{a_{103}}$$

OpenStudy (anonymous):

so it is 9 to 1 ? :)